Calculate the mass of aluminium needed to react with 50 g of cr2o3
1
Expert's answer
2020-11-04T14:22:53-0500

2Al + Cr2O3 → Al2O3 + 2Cr

"n(Cr2O3)=\frac{m}{M}=\frac{50}{151.989}=0.32897"

find n(Al):

"\frac{x}{2}=\frac{0.32897}{1}"

x=0.6579

"m=n\times M=0.6579\times26.9815=17.75" g